TCCON > Sites > East Trout Lake > Instrument History > Site Visit 2017-12-18 to 2017-12-21

Site Visit 2017-12-18 to 2017-12-21

To do:

  • Realign the IFS125HR
  • Install TriScroll instead of Iwata

December 18, 2017

  • Arrived at the lab around 5 pm. Waited for the bedtime story to finish. Put system in Hold and then SW Status Shutdown Quickly.
  • Took evacuated lamp scan and then vented scan with OPUS - looks like left shoulder is a little soft
  • Looked as laser modulation efficiency. MOPD MEs were higher than the ZPD MEs - indicates a shear misalignment. After adjusting the fixed cube corner (just the vertical adjustment screw a ~half turn clockwise), the MEs look better (though nothing as nice as Lamont).
  • This is the "before" adjustment measurement of ME:
  Vmax Vmin Pk-Pk Vzero ME
Front LASA -2.76 -7.68 4.92 -9 0.650793651
Front LASB -4.24 -8.16 3.92 -9.32 0.628205128
Back LASA -2.76 -7.76 5 -9 0.668449198
Back LASB -4.24 -8.16 3.92 -9.32 0.628205128
  • This is the "after" adjustment of ME: 
  Vmax Vmin Pk-Pk Vzero ME
Back LASA -2.72 -7.84 5.12 -9 0.688172043
Back LASB -4.2 -8.24 4.04 -9.32 0.651612903
Front LASA -2.64 -7.88 5.24 -9 0.700534759
Front LASB -4.12 -8.24 4.12 -9.32 0.656050955
  • Lamont has MEs above 0.8.
  • We noted that the lamp image was not centered on the input pinhole and that the lamp image was not centered on the exit pinhole.
  • We adjusted the lamp image on the input pinhole, but did not adjust the lamp image on the exit pinhole.
    • This is because we think this is better done on the sun (if we see any this week).
    • However, the continuum levels of the spectra recorded since the last alignment are lower than they were previously, so it's entirely possible that the two pinholes are not coaligned.
  • At 8:20 pm, starting coalignment with IR - flat mirror adjustment.
  • Confused about the isolation valve on the TriScroll - don't know whether it's needed or whether we can just put it in series with the other sensor and tube to the IFS.
  • At 8:45 pm, started vertical adjustment screw. Will do a couple of these, and then call it a night (and do some LINEFIT).
    • Noticed lines getting shallower; backed off from that direction and started the other direction.
  • At night, looked at linefit: 0.85 ME at MOPD, so we have more work to do. Will start by continuing in this direction and then backing off in steps of 5% instead of 10%.

Back to top

December 19, 2018

  • Arrived at the lab in darkness at 8 am.
  • Took scan without changing anything - looks roughly the same as last night's last scan
  • Continued in same direction as before at 10% increment.
  • Achieved what we thought was about the best we could do: ME at MOPD is 1.03 according to LINEFIT.
  • Noticed that the first mirror inside the IFS is rather dirty and has a strange circle pattern on it. Not sure why.
  • Aligned the solar beam into APT and AP1 was nicely centred.
  • Set up the TriScroll, which has an isolation valve.
  • We cleaned all the tubing as best we could of the black dust.

IMG_3686.JPG

  • On first pumpdown (with a tee connecting to the Micropirani and a blank), it achieved a vacuum of 0.04 hPa.
  • We then connected the tubing to the IFS, and it only achieved ~1.3 hPa. We disconnected all the tubing and cleaned it all as thoroughly as possible. There was still black dust inside the tubing.
  • After cleaning all the tubing, the ultimate pressure was 0.11 hPa, which is obviously better, but not 0.04 hPa.
  • Then we disconnecting the tubing and put the blank back in place. Now the ultimate pressure is ~0.06 hPa.
    • I hope we didn't do any damage to the new pump with the residual black dust.
  • We'll clean the connections around the isolation valve and add a bit of vacuum grease and try again.
  • The pump manual suggested running the pump on atmospheric pressure for a minute or two to flush out any dust or dirt. We did that as well, and it sounded a bit like it was struggling. We turned off the pump after 1 minute and put the blank back on. After 2 minutes, the pressure was 0.06 hPa. We'll wait a bit longer to see if it can reach 0.04 hPa. After 13 minutes, it reached 0.04 hPa.
    • Now we'll connect the tube to the IFS and evacuate it. We reached 0.04 hPa after a few minutes.
  • Pumped down IFS to 0.4 hPa without too much trouble.
  • Running cell scan with OPUS to see whether pumping down affected ILS. ILS is now 1.02 ME at MOPD 45 cm with APT=1300. We'll call it there.
  • Cleaned CaF2 window in ceiling - some dead flies and dust gently wiped off with dry Kimwipe.

Back to top

December 20, 2017

  • To do:
    • Peak detectors and take ghost spectra.
    • Set up alignment kit to suggest improvements and take pictures of current alignment (without touching optics!).
    • Take pictures of future location of Phenocam.
    • Show Joseph where input aperture will be mounted (to do in the summer).
  • Peaked detectors: required very little change. Gained <1% (from 2.044V to 2.057V) on InGaAs and basically nothing on InSb (11650 counts to 11670 counts according to OPUS check signal)
  • Looked at ghosts. Did not see any ghosts when XSM was switched on, as expected. Turning it off required editing a file, which we were not comfortable doing, and given that we didn't see any ghosts, it didn't seem worthwhile. Also noticed that the ghost filter I purchased was not the right one for the job (it let wavenumbers below 5400 cm-1 through). See the Technical Information section for details of what I should have bought.
  • Set up alignment kit. Have many suggested improvements.
  • Took pictures of the future location of the Phenocam. Will clearly not be upside down when installed...

IMG_3756.JPG

  • Showed Joseph where the input aperture will be mounted.
    • This is the largest aperture, but you'll have to check the signal levels. We ultimately want to be able to use 1 mm APT without saturating the InGaAs or InSb detectors near noon in the summer (you want MXY<0.9). You have two smaller apertures in the top drawer, and if you take out the current aperture, you have the largest possible size.

IMG_3716.JPG

  • The ILS under vacuum with APT=1000 at 75cm OPD looks good: ME=1.02.

 

Back to top

December 21, 2017

  • Pack up and return to YYZ.
  • Took the alignment kit and Iwata to Purolator and had them ship back to Toronto, but delayed until December 27 (to avoid arriving in Toronto when U of T is closed).

Back to top

Tag page

Files 3

FileSizeDateAttached by 
 et20171219lhffaa.020
No description
43.58 MB12:48, 10 Jul 2018mendoncaActions
 et20171219lhffaa.021
No description
43.58 MB12:48, 10 Jul 2018mendoncaActions
 et_HCL_45_1mm.xpm
ETL HCl xpm used to take call measurments
18.55 kB18:15, 20 Dec 2017mendoncaActions